You are going to make a box out of cardboard. The box needs to measure 10 cm. by 5 cm. by 5 cm. 10% of the final product will be wasted in the cutting process. How much cardboard do you need???

Thanks!!

Answer by checkley71(8403)

IF THE FINAL PRODUCT IS 10X5X5.
10*15=150
2*5*5=50
THIS TOTAL SURFACE AREA IS 100+50=200.
THE 10% WASTE IS 200*.1=20 SQUARE CM. THUS THE TOTAL AREA OF THE CARDBOARD WAS:
200+20=220 SQ CM.